OpenStudy (anonymous):

what is the implicit form of the function y = x^3 ?

OpenStudy (anonymous):

Since it is currently written in explicit form (since y=something with only x's) I would say that the implicit form would be either y^(1/3)=x or y-x^3=0, because both of those would eliminate the y=f(x) portion. Of those two, I can't say which is best--just that both are in an implicit form.